# Uncertainty-Calibrated Test-Time Model Adaptation without Forgetting

### v1

@@ Abstract
Out-of-Distribution Generalization, Test-Time Adaptation, Calibration, Catastrophic, Forgetting, Robustness.
@@ 1 Introduction
Over-Confident Predictions. Existing methods like Tent [12] and SAR [18] primarily focus on reducing the model’s prediction uncertainty through unsupervised prediction entropy minimization, which enhances the model’s predicting confidence, thereby learning a decision boundary positioned in low-density sample regions [19, 20]. However, a critical oversight of these methods is their neglect of the irreducible data uncertainty inherent in test samples. They encourage highly certain predictions for all test samples even in cases of underlying data uncertainty or even if the prediction is wrong, leading to overly certain (or overconfident) predictions. This phenomenon is particularly concerning in high-risk applications [21, 22], posing potential safety risks.
@@
To mitigate the over-confident issue, it is essential to differentiate between various origins of uncertainty in TTA, i.e., reducible model uncertainty and inherent data uncertainty. Inspired by the human tendency to base predictions on available knowledge and problem difficulty, rather than defaulting to high certainty in unfamiliar situations, we propose to utilize the model uncertainty for calibrated uncertainty reduction and leverage the inherent data uncertainty to directly re-calibrate predicting confidence.
@@
To this end, we further propose EATA with Calibration, namely EATA-C. To be specific, EATA-C estimates model uncertainty by measuring the divergence between predictions from the full network and its randomly sampled sub-networks. This allows for the reduction of model uncertainty by minimizing the divergence, promoting consistent rather than overconfident predictions during testing. Additionally, we develop a simple but effective indicator of data uncertainty based on the disagreement among predicted labels, which suggests whether a sample nears decision boundaries. Then, we devise an auxiliary min-max entropy regularizer to selectively lower confidence for inherently uncertain samples through an entropy maximization loss, while simultaneously boosting confidence for certain samples via an entropy minimization loss. EATA-C differs from EATA on the adaptation objective, while it still benefits from the active sample selection criterion and anti-forgetting Fisher regularization proposed in EATA. We summarize our main contributions as follows.
@@
  • We propose an Efficient Anti-forgetting Test-time Adaptation (EATA) method. Specifically, we develop an active sample identification scheme to filter out non-reliable and redundant test samples from model adaptation, thereby improving adaptation efficiency by reducing backward propagation. Moreover, we extend the label-dependent Fisher regularizer to test samples with pseudo label generation, which prevents drastic changes in important model weights and thus alleviates the issue of model forgetting on in-distribution test samples.
@@
  • We further propose EATA with Calibration (EATA-C), which decouples the model and data uncertainty to achieve confidence-calibrated TTA. Specifically, EATA-C estimates model uncertainty using the divergence between full network and sub-network predictions, incorporating a consistency loss to reduce this uncertainty without leading to overconfidence. For data uncertainty, EATA-C utilizes label prediction disagreements, employing min-max entropy regularization to selectively adjust prediction confidence for effective re-calibration.
@@
  • We demonstrate that our proposed EATA method improves both the performance and efficiency of test-time adaptation and also alleviates the long-neglected catastrophic forgetting issue, and EATA-C further achieves better performance and calibration without additional backward computations compared with EATA.
@@ 2 Related Work
Test-Time Adaptation (TTA) aims to improve model accuracy on OOD test data through model adaptation with test samples, which has been successfully applied to various tasks, such as image classification [11, 12], image super-resolution [24], image captioning [25], object detection [26], video recognition [27, 28], etc. Existing test-time training methods, e.g., TTT [11], TTT++ [13], TTT-MAE [29], and MT3 [30], jointly train a source model via both supervised and self-supervised objectives, and then adapt the model via self-supervised objective at test time. This pipeline, however, necessitates both self-supervised head and test data in adaptation, while training such self-supervised head can be computation-consuming [29]. To address this, some methods have been proposed to adapt a model with only test data, including batchnorm statistics adaptation [31, 32, 33], prediction consistency maximization over different augmentations [34], and classifier adjustment [35]. Specifically, Tent [12] updates the model to minimize the entropy of predictions at test time. MEMO [14] further augments test samples for marginal entropy minimization to enhance robustness. Our work also alleviates the dependency on self-supervision heads and seeks to address the key limitations of prior works (i.e., efficiency hurdle, catastrophic forgetting, and overconfidence) to make TTA more practical in real-world applications.
@@ 3 Problem Formulation
Moreover, existing entropy-based test-time adaptation methods like Tent [12] and SAR [18] consistently encourage producing one-hot predictions with high confidence. However, in practice, incoming test samples can be severely corrupted [9], inheriting irreducible data uncertainty. Considering the data uncertainty, these test samples ideally should be predicted with relatively low confidence. Nevertheless, the data uncertainty is often overlooked by methods based on entropy minimization. Consequently, the adapted model often produces highly confident predictions (called overconfidence), even when the test samples are uncertain. Such misleading predictions raise potential safety concerns for real-world application scenarios. We empirically demonstrate the issue of overconfidence in Figure 7(a) and Table II.
@@ 4.1 Sample Efficient Entropy Minimization
Figure 2: Effect of different test samples in test-time entropy minimization [12]. We adapt a model on partial samples (top $p$samples with the highest or lowest entropy values), and then evaluate the adapted model on all test samples. Results are obtained on ImageNet-C (Gaussian noise, level 3) and ResNet-50 (base accuracy is 27.6%). Introducing more samples with high entropy values into adaptation will hurt the adaptation performance.
@@ 4.3 Consistency-Based Uncertainty Minimization
As mentioned in Section 4.1, EATA enhances model performance by minimizing prediction entropy. This strategy aims to reduce uncertainty in predictions and learn a decision boundary in the low-density regions of the test samples [58, 59]. However, a persistent challenge with entropy minimization is its tendency to yield overly certain predictions, even in cases where the test samples themselves are inherently uncertain, such as severely corrupted images [9]. Thus, EATA may still result in overconfident predictions that fail to accurately account for the inherent data uncertainty. To address this, we further propose EATA with Calibration (EATA-C), shall be depicted in Sections 4.3 and 4.4.
In EATA-C, we first propose a consistency-based loss to quantify and optimize the model uncertainty. Our method is inspired by MC Dropout [60], which has shown promising performance in estimating the model uncertainty through the divergence of multiple dropout-enabled predictions. In our context, considering adaptation efficiency, we define the model uncertainty as the KL divergence [61] between the full network prediction and randomly sampled sub-network prediction. Here, we use only two predictions and the former is indispensable since we select it as the final prediction. Thus, at test time, we minimize divergence to encourage consistent predictions instead of confidence-related entropy loss, thereby alleviating the overconfidence issue.
@@
where $D_{KL}(\cdot||\cdot)$ denotes Kullback-Leibler divergence [61] and $p$ is a constant for balancing. Here, we calculate the divergence between $\hat{{\bf y}}$ and $\hat{{\bf y}}_{fuse}$ (rather than $\hat{{\bf y}}_{sub}$), since encouraging the sub-network to achieve the same performance as the full one is relatively hard. Thus, inspired by label smoothing [62], we softly fuse $\hat{{\bf y}}$ and $\hat{{\bf y}}_{sub}$ in Eqn. (13) for divergence optimization. Note that, during optimization, we conduct a unidirectional alignment from the sub-network to the full network, as the full network typically exhibits stronger generalization capabilities. To this end, we detach the gradient from $\hat{{\bf y}}$ and concentrate the optimization solely on $\hat{{\bf y}}_{sub}$. This strategy is designed to facilitate knowledge transfer from the full network to its sub-network, thereby enhancing the sub-network’s performance while simultaneously reducing the model uncertainty of the full network.
Remark on Efficiency. Although consistency loss requires two forward passes from both the full network and the sub-network for each sample, the full network’s forward pass is gradient-free without back-propagation and the sub-network’s forward/backward pass is less computationally intensive. Moreover, we only perform sub-network’s forward/backward pass on the selected reliable and non-redundant samples as outlined in Algorithm 1. As a result, the use of consistency loss remains efficient.
@@ 4.4 Calibrated Min-Max Entropy Regularization
In this section, we re-calibrate the model’s prediction uncertainty in a manner that is sensitive to individual samples. This process involves categorizing samples into two distinct groups—‘certain’ and ‘uncertain’—based on the aforementioned prediction consistency. This identification is achieved by comparing the predicted labels between the full network and a sub-network. Samples exhibiting mismatched predictions are deemed ‘uncertain’, indicating either high intrinsic data uncertainty or a high level of complexity that challenges the model’s recognition capability. For these uncertain samples, we aim to lower their predictive confidence. This is achieved by maximizing the prediction entropy, effectively acknowledging the model’s lack of confidence in these cases. Conversely, for samples where predictions are consistent, labeled as ‘certain’, we conduct the opposite strategy, i.e., boosting prediction confidence through entropy minimization. Formally, this min-max entropy regularization optimization problem is defined by:
@@
$C({\bf x})=\begin{cases}1,&\text{if $\operatorname*{arg\,max}(\hat{{\bf y}})=\operatorname*{arg\,max}(\hat{{\bf y}}_{sub})$,}\\ -1,&\text{if $\operatorname*{arg\,max}(\hat{{\bf y}})\neq\operatorname*{arg\,max}(\hat{{\bf y}}_{sub})$,}\\ \end{cases}$ (15)
@@
Overall Objective. The methods proposed in Sections 4.3 and 4.4 are devised to address the overconfident issue in TTA, but still suffer from catastrophic forgetting when the important model weights for the in-distribution domain are modified significantly during adaptation. Therefore, we jointly optimize the model with the anti-forgetting regularizer and further reduce the required backward computations with the active sample selection criterion proposed in EATA. Then, the overall objective of EATA-C can be formulated as:
@@ 5 Experiments
TABLE II: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, normal-$\uparrow$) and Expected Calibration Error(%,normal-$\downarrow$). “BN” and “LN” denote batch and layer normalization, respectively. The bold number indicates the best result and the underlined number indicates the second best result. All results are evaluated under the lifelong adaptation scenario except for Tent [12] and MEMO [14], which suffer severely from error accumulation. We use * and ${}^{\dagger}$ to denote episodic and single-domain adaptation, respectively.
@@
EATA (Ours) Acc. 35.6 38.7 37.5 35.9 36.1 47.6 53.1 50.6 45.6 59.5 67.1 45.2 57.6 59.8 55.4 48.36
ECE 10.5 13.4 14.7 18.5 18.5 14.7 12.8 14.1 16.1 11.6 10.4 18.6 13.3 13.1 13.9 14.28 50,000 29,721
EATA-C (Ours) Acc. 37.2 40.9 39.9 36.6 37.1 48.5 52.7 51.9 47.2 60.4 67.0 49.0 57.6 60.1 56.1 49.48
ECE 7.1 6.7 7.2 10.1 9.5 5.9 4.9 4.8 5.8 3.2 3.2 6.0 4.0 3.9 3.9 5.74 83,312 33,312
@@
EATA (Ours) Acc. 50.5 55.6 56.0 54.9 56.3 61.1 59.8 64.3 64.0 70.1 77.4 65.5 63.1 70.4 69.5 62.57
ECE 10.6 13.9 15.7 16.4 17.3 15.9 17.5 14.8 15.7 12.8 9.0 15.4 17.5 13.3 13.9 14.63 50,000 36,688
EATA-C (Ours) Acc. 56.8 60.2 59.8 58.0 60.9 65.2 65.5 69.7 67.8 74.0 78.9 66.7 70.3 74.1 71.8 66.65
ECE 5.2 5.1 5.9 7.1 6.2 5.6 5.9 5.0 5.0 4.1 3.3 5.4 4.7 4.2 4.7 5.14 83,184 33,184
@@
Model Acc. (%) ECE (%) #Forwards #Backwards

### v2

@@ Abstract
Out-of-Distribution Generalization, Test-Time Adaptation, Confidence Calibration, Catastrophic Forgetting.
@@ 1 Introduction
Over-Confident Predictions. Existing methods like Tent [12] and SAR [18] primarily rely on test-time entropy minimization for model adaptation, which greedily enhances the model’s confidence and minimizes the predictive uncertainty for test samples, without distinguishing between model-induced and data-induced uncertainties. Consequently, even when the input data is naturally complex or highly corrupted (i.e., with irreducible data uncertainty), the model is forced to make one-hot confident predictions where it should remain uncertain, leading to over-confident and potentially incorrect outputs. This phenomenon is particularly concerning in high-risk applications, such as autonomous driving [19] and medical diagnosis [20], posing potential safety risks.
@@
To mitigate overconfidence, we differentiate between various origins of uncertainty in TTA: 1) Reducible model uncertainty, which arises from not knowing the optimal model parameters to describe the data due to insufficient training [22]; 2) Irreducible data uncertainty that arises from inherent noise or variability in the data, cannot be reduced by additional training [23]. Based on their characteristics, we aim to reduce the model uncertainty at test time for domain adaptation, while accurately reflecting data uncertainty in model predictions to ensure confidence calibration.
@@
To this end, we further devise EATA with Calibration, namely EATA-C. Specifically, EATA-C estimates model uncertainty by measuring the divergence between predictions from the full network and its randomly sampled sub-networks. By minimizing this divergence, our EATA-C reduces model uncertainty and promotes consistent, rather than overconfident, predictions for model adaptation during testing. Additionally, we introduce a data uncertainty indicator based on prediction disagreement, which effectively detects ambiguous samples near decision boundaries where conflicting predictions are more likely to occur. We then incorporate a min-max entropy regularizer to selectively adjust the prediction confidence based on this data uncertainty estimation. Note that EATA-C differs from EATA on the adaptation objective, while it still benefits from the active sample selection criterion and anti-forgetting Fisher regularization proposed in EATA. We summarize our main contributions as follows.
@@
  • We propose an Efficient Anti-forgetting Test-time Adaptation (EATA) method. Specifically, we reveal that test samples contribute differently to adaptation, and develop an active sample identification scheme to filter out non-reliable and redundant test samples from adaptation, thereby improving TTA efficiency. Moreover, we extend the label-dependent Fisher regularizer to test samples with pseudo label generation, which prevents drastic changes in important model weights and helps alleviate the issue of model forgetting on in-distribution test samples.
@@
  • We further introduce EATA with Calibration (EATA-C), which differentiates between reducible and irreducible uncertainty during testing to design a calibration-driven learning objective. Specifically, EATA-C estimates model uncertainty using the divergence between full network and sub-network predictions, incorporating a consistency loss to reduce this uncertainty for adaptation. Regarding data uncertainty, EATA-C leverages prediction disagreements and applies min-max entropy regularization to selectively adjust confidence for calibration enhancement.
@@
  • We demonstrate that our proposed EATA method improves both the performance and efficiency of test-time adaptation and also alleviates the long-neglected catastrophic forgetting issue. Our EATA-C further achieves better performance and calibration, with computational and memory efficiency comparable to EATA.
A short version of this work was published in ICML 2022 [24]. This paper extends our preliminary version from the following aspects: 1) We explore the calibrated test-time adaptation, which aims to provide calibrated predicting confidence that reflects the true likelihood of correctness during unsupervised adaptation; 2) To solve the over-confident issue, we develop a test-time consistency loss that leverages the reducible model uncertainty for calibrated uncertainty reduction, and devise a min-max entropy regularizer to re-calibrate predicting confidence based on the inherent data uncertainty; 3) We provide analyses about the impact of different uncertainty reduction strategies, empirically verifying that our consistency loss overcomes the issue of over-fitting and over-confident in entropy minimization loss when adapting to the test data; 4) We provide extensive new empirical evaluations on image classification and semantic segmentation tasks with various model architectures, demonstrating that EATA-C achieves substantially better performance and calibration over EATA, e.g., improving accuracy by 6.5%, while reducing calibration error by relatively 64.9% on ImageNet-C dataset with ViT-Base [25].
@@ 2 Related Work
Test-Time Adaptation (TTA) aims to improve model accuracy on OOD test data through model adaptation with test samples. Existing test-time training methods, e.g., TTT [11], TTT++ [13], TTT-MAE [26], and MT3 [27], jointly train a source model via both supervised and self-supervised objectives, and then adapt the model via self-supervised objective at test time. This pipeline, however, necessitates both self-supervised head and test data in adaptation, while training such self-supervised head can be computation-consuming [26]. To address this, some methods have been proposed to adapt a model with only test data, including batchnorm statistics adaptation [28, 29, 30], prediction consistency maximization over different augmentations [31], and classifier adjustment [32]. Specifically, Tent [12] updates the model to minimize the entropy of predictions at test time. MEMO [14] further augments test samples for marginal entropy minimization to enhance robustness. Our work also alleviates the dependency on self-supervision heads and seeks to address the key limitations of prior works (i.e., efficiency hurdle, catastrophic forgetting, and overconfidence) to make TTA more practical in real-world applications.
@@ 3 Problem Formulation
Moreover, existing entropy-based test-time adaptation methods like Tent [12] and SAR [18] consistently encourage the model to produce one-hot highly confident predictions. However, in practice, input test samples can be naturally complex and severely corrupted [9], resulting in irreducible data uncertainty. Ideally, these samples should be predicted with relatively low confidence to reflect their ambiguity. Nevertheless, the data uncertainty is often overlooked by methods based on entropy minimization,causing the adapted model to produce highly confident predictions (called overconfidence), even when predictions should remain uncertain. Such misleading predictions raise potential safety concerns for real-world application scenarios. We empirically demonstrate the issue of overconfidence in Figure 11(a) and Table II.
@@ 4.1 Sample Efficient Entropy Minimization
Figure 2: Effect of different test samples in test-time entropy minimization [12]. We adapt a model on partial samples (top pp% samples with the highest or lowest entropy values), and then evaluate the adapted model on all test samples. Results are obtained on ImageNet-C (Gaussian noise, level 3) and ResNet-50 (base accuracy is 27.6%). Introducing more samples with high entropy values into adaptation will hurt the adaptation performance.
@@ 4.3 Consistency-Based Uncertainty Minimization
As mentioned in Section 4.1, EATA conducts model adaptation by prediction entropy minimization. This strategy aims to reduce uncertainty in predictions and learn decision boundaries in low-density regions of the test samples [55, 56]. However, a persistent limitation of entropy minimization is its tendency to yield overly certain predictions, where the model is forced to make one-hot confident predictions even for ambiguous input data with irreducible data uncertainty. Thus, EATA may still result in overconfident predictions that do not accurately reflect the inherent data uncertainty. To address this, we further propose EATA with Calibration (EATA-C), shall be depicted in Sections 4.3 and 4.4.
In EATA-C, we first propose a consistency-based loss to quantify and optimize the model uncertainty. Our method is inspired by MC Dropout [57], which has shown promising performance in estimating the model uncertainty through the divergence of multiple dropout-enabled predictions. In our context, considering adaptation efficiency, we define the model uncertainty as the KL divergence [58] between the full network prediction and randomly sampled sub-network prediction. Here, we use only two predictions and the former is indispensable since we select it as the final prediction. During TTA, we minimize this divergence to promote consistent predictions for model updates, rather than greedily increasing confidence which can result in overconfident outputs.
@@
where $D_{KL}(\cdot||\cdot)$ denotes Kullback-Leibler divergence [58] and $p$ is a constant for smoothing. Here, we calculate the divergence between $\hat{{\bf y}}_{sub}$ and $\hat{{\bf y}}_{fuse}$ (rather than $\hat{{\bf y}}$), since encouraging the sub-network to achieve the same performance as the full one is relatively hard. Thus, inspired by label smoothing [59], we softly fuse $\hat{{\bf y}}$ and $\hat{{\bf y}}_{sub}$ in Eqn. (13) for divergence optimization. Note that, during optimization, we conduct a unidirectional alignment from the sub-network to the full network, as the full network typically exhibits stronger generalization capabilities. To this end, we detach the gradient from $\hat{{\bf y}}$ and concentrate the optimization solely on $\hat{{\bf y}}_{sub}$. This strategy is designed to facilitate knowledge transfer from the full network to its sub-network, thereby enhancing the sub-network’s performance while reducing the full networks’ model uncertainty to adapt the network to the test domain.
Remark on Efficiency. Although consistency loss requires two forward passes from both the full network and the sub-network for each sample, the full network’s forward pass is gradient-free without back-propagation and the sub-network’s forward/backward pass is less computationally intensive. Moreover, we only perform sub-network’s forward/backward passes on the selected reliable and non-redundant samples as outlined in Algorithm 1. As a result, the use of consistency loss remains efficient as shown in Table VII.
@@ 4.4 Calibrated Min-Max Entropy Regularization
In this section, we re-calibrate the model’s prediction uncertainty in a manner that is sensitive to individual samples. This process involves categorizing samples into two distinct groups—‘certain’ and ‘uncertain’—based on the aforementioned prediction consistency. This design is inspired by margin-based learning approaches [61, 62] which indicated that samples near decision boundaries are inherently more uncertain and have been well justified with theoretical guarantees. Specifically, we achieve categorization by comparing the predicted labels between the full network and a sub-network. Samples that exhibit mismatched predictions are deemed ‘uncertain’, suggesting their proximity to decision boundaries and high intrinsic data uncertainty. Note that unlike the consistency loss that measures model uncertainty, in which samples across the data space may yield low consistency loss, data uncertainty is reflected more prominently through prediction disagreements near the decision boundary (see Figure C for illustration). For identified uncertain samples, we aim to lower their predictive confidence by maximizing the prediction entropy, effectively acknowledging the model’s lack of confidence in these cases. Conversely, for samples where predictions are consistent, labeled as ‘certain’, we conduct the opposite strategy, i.e., boosting prediction confidence through entropy minimization. Formally, this min-max entropy regularization optimization problem is defined by:
@@
[…]
@@
Overall Objective. The methods proposed in Sections 4.3 and 4.4 are devised to address the overconfident issue in TTA, but still suffer from catastrophic forgetting when important model weights for the in-distribution domain are significantly modified during adaptation. Therefore, we jointly optimize the model with the anti-forgetting regularizer and further reduce the required backward computations with the active sample selection criterion in EATA. Then, the overall objective of EATA-C can be formulated as:
@@ 5 Experiments
TABLE II: Comparison with state-of-the-art methods on ImageNet-C with the highest severity level 5 regarding Corruption Accuracy(%, $\uparrow$) and Expected Calibration Error($\downarrow$). “BN” and “LN” denote batch and layer normalization, respectively. The bold number indicates the best result and the underlined number indicates the second best result. All results are evaluated under the lifelong adaptation scenario except for Tent [12] and MEMO [14], which suffer severely from error accumulation. We use * and to denote episodic and single-domain adaptation, respectively.
@@
ROID Acc. 36.7 38.6 35.9 29.1 28.8 40.6 46.5 49.5 41.8 55.6 65.3 43.5 53.6 56.1 52.8 44.95 80,380 80,380
ECE 36.6 38.5 35.8 29.0 28.7 40.5 46.4 49.4 41.7 55.5 65.2 43.4 53.5 56.0 52.7 44.85
RDump Acc. 35.6 34.4 36.1 32.7 34.6 45.5 51.6 50.3 44.1 59.9 66.8 46.3 56.8 59.1 54.7 47.23 50,000 26,375
ECE 10.5 11.4 12.6 13.9 14.4 9.4 8.1 8.1 9.4 6.3 5.5 11.8 6.1 7.0 6.3 9.40
[…]
ECE 7.9 5.4 2.4 2.0 1.8 1.9 3.8 19.5 67.6 76.4 86.2 91.0 90.4 91.3 91.3 42.60
Acc. 35.6 38.7 37.5 35.9 36.1 47.6 53.1 50.6 45.6 59.5 67.1 45.2 57.6 59.8 55.4 48.36
EATA (Ours) ECE 10.5 13.4 14.7 18.5 18.5 14.7 12.8 14.1 16.1 11.6 10.4 18.6 13.3 13.1 13.9 14.28 50,000 29,721
Acc. 37.2 40.9 39.9 36.6 37.1 48.5 52.7 51.9 47.2 60.4 67.0 49.0 57.6 60.1 56.1 49.48
EATA-C (Ours) ECE 7.1 6.7 7.2 10.1 9.5 5.9 4.9 4.8 5.8 3.2 3.2 6.0 4.0 3.9 3.9 5.74 83,312 33,312
@@
ROID Acc. 48.8 49.5 49.0 54.1 54.2 58.8 55.7 62.7 61.3 69.7 77.0 65.5 64.1 69.6 68.3 60.56 80,739 80,739
ECE 48.7 49.4 48.9 54.0 54.1 58.7 55.6 62.6 61.2 69.6 76.9 65.4 64.0 69.5 68.2 60.46
RDump Acc. 50.5 48.4 51.8 54.0 55.3 59.4 55.9 63.3 60.7 70.8 76.8 66.8 60.9 69.5 68.4 60.84 50,000 32,050
ECE 10.6 9.7 10.5 8.9 9.4 7.9 10.0 6.9 7.4 5.4 3.2 6.7 7.6 5.1 5.0 7.62
[…]
ECE 10.6 13.7 13.7 14.5 14.4 11.3 11.6 9.4 10.2 7.8 3.9 10.5 9.8 7.3 7.4 10.40
Acc. 50.5 55.6 56.0 54.9 56.3 61.1 59.8 64.3 64.0 70.1 77.4 65.5 63.1 70.4 69.5 62.57
EATA (Ours) ECE 10.6 13.9 15.7 16.4 17.3 15.9 17.5 14.8 15.7 12.8 9.0 15.4 17.5 13.3 13.9 14.63 50,000 36,688
Acc. 56.8 60.2 59.8 58.0 60.9 65.2 65.5 69.7 67.8 74.0 78.9 66.7 70.3 74.1 71.8 66.65
EATA-C (Ours) ECE 5.2 5.1 5.9 7.1 6.2 5.6 5.9 5.0 5.0 4.1 3.3 5.4 4.7 4.2 4.7 5.14 83,184 33,184
@@
Model
Acc. (%)
ECE (%)
#Forwards
#Backwards
@@
$\bullet$~BN [29] 40.4(+2.4) 13.4(-4.3) 30,000 0
$\bullet$~Tent [12] 42.3(+4.3) 17.8(+0.1) 30,000 30,000
[…]
$\bullet$~CoTTA [63] 42.4(+4.4) 15.8(-1.9) 90,000 30,000
$\bullet$~SAR [18] 42.7(+4.7) 14.6(-3.1) 47,755 32,877
$\bullet$~ROID [64] 48.6(+10.5) 48.1(+30.4) 48,303 48,303
[…]
$\bullet$~EATA (Ours) 44.9(+6.9) 16.7(-1.0) 30,000 5,417
$\bullet$~EATA-C (Ours) 47.1(+9.1) 13.3(-4.4) 35,122 5,122
ViT(LN) 52.5 5.0 30,000 0
$\bullet$~Tent [12] 54.2(+1.7) 7.4(+2.4) 30,000 30,000
[…]
$\bullet$~CoTTA [63] 56.4(+3.9) 7.4(+2.4) 90,000 30,000
$\bullet$~SAR [18] 55.0(+2.5) 5.2(+0.2) 47,119 33,844
$\bullet$~ROID [64] 62.2(+9.7) 61.7(+56.7) 49,795 49,795
[…]
$\bullet$~EATA (Ours) 58.2(+5.7) 5.8(+0.8) 30,000 6,053
$\bullet$~EATA-C (Ours) 64.2(+11.7) 3.9(-1.1) 36,395 6,395
Datasets and Models. We conduct experiments on three benchmark datasets for OOD generalization: ImageNet-C [9] (contains corrupted images in 15 types of 4 main categories and each type has 5 severity levels) and ImageNet-R [66] for image classification; and ACDC [67] for semantic segmentation. We use ResNet-50 (R-50) [1] and ViT-Base (ViT) [25] for ImageNet experiments, and Segformer-B5 [68] for ACDC [67] experiments. The models are trained on ImageNet or CityScapes [69] training set with stochastic depth regularization [60] and tested on clean or OOD test sets.
Compared Methods. We compare with the following state-of-the-art methods. BN adaptation [29] updates batch normalization statistics on test samples. Tent [12] minimizes the entropy of test samples during testing. MEMO [14] maximizes the prediction consistency of different augmented copies regarding a given test sample. SAR [18] selects reliable samples for test time sharpness-aware entropy minimization. CoTTA [63] and DAT [70] minimize the cross entropy between the student network and its mean teacher during testing. RDump [71] periodically resets model parameters based on our EATA. TEA [65] employs energy-based contrastive learning with negative sample generation. ROID [64] minimizes the diversity-weighted soft likelihood ratio loss. We denote EATA without weight regularization in Eqn. (9) as efficient test-time adaptation (ETA). More ablative methods can be found in Table VI.
Round 1 4 7 10 All
BN Stats Adapt 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 62.3 38.0 54.6 53.0 52.0(-4.7)
Tent (lifelong) 69.0 40.2 60.0 57.3 66.6 36.6 58.9 54.2 64.6 33.4 55.9 51.6 62.5 30.4 52.6 48.7 52.7(-4.0)
CoTTA 70.9 41.1 62.4 59.7 70.8 40.6 62.6 59.7 70.8 40.5 62.6 59.7 70.8 40.5 62.6 59.7 58.4(+1.7)
DAT 71.7 44.6 63.8 62.2 68.0 42.0 60.9 59.4 66.1 40.6 59.8 57.8 63.8 39.6 58.2 55.4 57.0(+0.3)
EATA 69.1 40.5 59.8 58.1 69.3 41.8 60.1 58.6 68.8 42.5 59.4 57.9 67.9 42.8 57.7 56.3 57.0(+0.3)
EATA-C 71.0 44.3 63.1 61.1 72.0 47.3 64.9 63.8 71.8 48.2 64.2 64.2 72.0 48.7 64.3 64.1 61.6(+4.9)
[…]
Results on ImageNet-C. From Table II, our EATA and EATA-C consistently surpass existing approaches regarding adaptation accuracy, e.g., the average accuracy of 48.4% (EATA) vs. 45.0% (ROID) on ResNet-50. Importantly, EATA yields a remarkable performance gain over its counterpart Tent, e.g., $33.4\%\rightarrow 50.5\%$ on Gaussian Noise with ViT-Base, suggesting the significance of removing samples with unreliable gradients and tackling samples differently in the TTA process. Our enhanced method, EATA-C, further boosts adaptation accuracy by a large margin, which consistently outperforms TEA and ROID in all 15 corruption types over both ResNet-50 and ViT-Base, suggesting our effectiveness. Note that besides achieving strong OOD performance, EATA also alleviates the forgetting on ID samples (see Figure 3), showing the effectiveness of our anti-forgetting regularization without limiting the learning ability during adaptation (see also Table VI for ablation).
[…]
Regarding calibration, existing methods consistently exhibit substantial calibration error (e.g., ROID and CoTTA are 60.46% and 27.59% on ECE with ViT-Base), suggesting miscalibration as a prevalent issue in the unsupervised test-time adaptation. By filtering unreliable samples to reduce noisy learning signals, EATA improves calibration over Tent (e.g., $11.7\%\rightarrow 10.5\%$ on Gaussian Noise with ResNet-50), though miscalibration is yet significant. By further decreasing reducible model uncertainty and reflecting data uncertainty in model predictions, our enhanced method, EATA-C, reduces the ECE of EATA by relatively 59.8% on ResNet-50 and 64.9% on ViT-Base, demonstrating the strong calibration effect of EATA-C across diverse datasets and architectures. In summary, while EATA-C improves performance and efficiency over the state of the art, our EATA-C further achieves high accuracy, well-calibrated prediction, and efficient computation simultaneously, establishing a new benchmark for test-time adaptation.
[…]
Results on CityScapes-to-ACDC. Following [63], we evaluate our method on the semantic segmentation task in a lifelong adaptation scenario. From Table IV, while DAT initially achieves higher mIOU, it tends to overfit, leading to significant performance degradation in subsequent adaptations. In contrast, our EATA maintains a more stable performance compared to DAT and Tent, by filtering unreliable predictions and preventing drastic changes to important model parameters. Moreover, by replacing entropy minimization with our consistency maximization objective for more robust learning signals, EATA-C achieves state-of-the-art performance, surpassing EATA by 4.6% and CoTTA by 3.2% in mIoU over ten adaptation rounds. More critically, our EATA-C showcases consistent improvement over lifelong adaptation, increasing the average mIOU on four datasets from 59.8% (first round) to 62.3% (tenth round), further highlighting our long-term effectiveness.
5.2 Demonstration of Preventing Forgetting
5.3 Ablation Studies
[…]
Effect of Components in EATA-C. Our EATA-C aims to achieve a favorable balance between accuracy, calibration, and efficiency. We conduct an ablation study to verify the effectiveness of each module as in Table VI. The results indicate the following findings: 1) Consistency Loss: Incorporating the consistency loss alone substantially enhances the source model’s robustness and reduces ECE; 2) Entropy Regularization: The min-max entropy regularizer further calibrates prediction confidence and leads to a slight improvement in accuracy, e.g., accuracy increases from 48.8% (Exp 10) to 49.0% (EATA-C), and ECE decreases from 5.1% to 4.6%; 3) Fisher Regularization: This anti-forgetting regularizer contributes to TTA stability, as in the lifelong TTA of Table II and Figure 3. In single-domain TTA, it also positively affects both ECE and accuracy, e.g., ECE decreases from 5.4% (ETA-C) to 4.6% (EATA-C) and accuracy improves from 48.9% to 49.0%; 4) Active Sample Selection: By filtering out unreliable and redundant test samples, active sample selection significantly boosts computational efficiency while maintaining or improving accuracy, e.g., accuracy increases from 48.5% (Exp 6) to 49.0% (EATA-C) while reducing the required backward passes by 35%. More discussions on wall-clock time and memory usage are provided in Table VII. These results collectively underscore the effectiveness of each component.
Figure 5: Effect of different similarity threshold $\epsilon$ in Eqn.( 7). Results obtained on ImageNet-C(Gaussian, level 5) with ResNet-50.
Figure 7: Effect of different balancing factor $\alpha$ in Eqn. (10). Results obtained on ImageNet-C(Gaussian, level 5) with ViT-Base.
Figure 8: Effect of different balancing factor $\beta$ in Eqn. (10). Results obtained on ImageNet-C(Gaussian, level 5) with ViT-Base.
Figure 9: Effect of different stochastic depth ratios in EATA-C. Results obtained on ImageNet-C(Gaussian, level 5) with ViT-Base.
Method Level 3 Level 5
Acc. (%) #Backwards Acc. (%) #Backwards
Baseline ($S({\bf x})\small{=}1$) 68.8 50,000 33.4 50,000
$S^{ent}{({\bf x})}$ (Eqn. 3) 70.7 45,302 49.6 37,636
$S{({\bf x})}$ (Eqn. 8) 70.8 36,057 50.4 28,168
[…]
TABLE VI: Effects of components in EATA-C. Results obtained on 15 corruptions of ImageNet-C (level 5) with ResNet-50 in single-domain TTA scenario, i.e., the model parameters are reset before adapting to a new domain. CL denotes consistency loss. ER denotes min-max entropy regularizer. FR denotes Fisher regularizer. $S$ denotes active sample selection.
Experiment CL ER FR $S$ Average
Acc. ECE #Forwards #Backwards
Source 39.8 7.5 50,000 0